# Causality for Earth Science - A Review on Time-series and Spatiotemporal Causality Methods

### v1

@@ 4.6.1 Marginal Structural Models
Marginal Structural Models (MSMs) (Robins et al., 2000) are a class of causal inference methods designed to estimate the causal effects of time-varying treatments in the presence of time-dependent confounding. While traditional methods like ordinary regression models may not properly handle time-varying treatments and confounders, MSMs were developed to address biases that can arise when using these traditional regression models to estimate causal effects in situations where treatments change over time. MSMs provide a framework to model and adjust for the dynamic nature of treatments, confounders, and their interdependencies using the IPTW weights. MSMs first employ IPTW to re-weight the data and emulate a hypothetical time-fixed treatment scenario. This mitigates confounding by making treated and untreated groups comparable. A weighted regression model is then employed to estimate the treatment effect, accounting for the dynamic confounding. MSMs rely on the assumptions of positivity and no unmeasured confounders. Though MSMs have become a cornerstone in time-varying causal inference in the fields of epidemiology (VanderWeele, 2009; Hernán et al., 2000), public health (Williamson and Ravani, 2017; VanderWeele et al., 2011) and social sciences(Bacak and Kennedy, 2015), they also have some limitations. Complex interactions between time-varying treatments and confounders, longitudinal missing data and information censoring can be challenging to model using MSM technique.
@@ 4.6.2 Convergent Cross Mapping
Convergent Cross Mapping (CCM) (Ye et al., 2015) is a nonlinear causal inference technique used to detect the presence of causal relationships between variables in time series data. It is particularly useful when the relationship between variables is complex and non-linear. CCM is based on the concept of time delay embedding, where time series data is transformed into higher-dimensional space by embedding time-delayed copies of the series.
@@ 4.6.4 Deep Representation Learning based Models
Causal inference methods based on representation learning or deep learning techniques (Bengio et al., 2013) learn the representation of input data by extracting features from the covariate space (Koch et al., 2021), where majority of the existing deep learning based methods are developed for i.i.d data (Koch et al., 2021). In these deep learning based CI methods, a single neural network (also called meta learner) can be trained to make predictions for both treatment and control groups individually to estimate the average treatment effect (ATE).
For time-series causal inference, researchers have proposed methodologies based on machine learning and deep learning models that also tackle the problem of time-varying confounding (Moraffah et al., 2021). Recurrent Marginal Structural Networks (R-MSN) (Lim, 2018) and Counterfactual Recurrent Network (CRN) (Bica et al., 2020b) are some of the recent models that claim to estimate causal effects in the presence of time-varying confounders, however, contrary to the claim, these methods are healthcare-specific and cannot be generalized for other domain areas like Earth science because these models require one-hot encoded treatment flags with multivariate combined dosage. Time Series Deconfounder - a multi-task method, leverages the assignment of multiple binary treatments over time to enable the estimation of treatment effects in the presence of multi-cause hidden confounders (Bica et al., 2020a). Taking this one step further, G-Net is a recently proposed method for time-varying dynamic treatment effect estimation (Li et al., 2021b). The method provides a recurrent-neural network based g-computation technique to estimate the propensity scores for handling time-varying confoundedness. Most recently, (Ali et al., 2023) proposed TCINet, a deep learning based counterfactual prediction model that leverages stabilized weighting (instead of IPTW weights) using probabilistic modeling. The authors evaluated TCINet on climate data and presented their findings on the relationship between Greenland blocking and summer sea ice melt within Barents Sea and Kara Sea.
Though deep representation learning methods are capable of automatically learning the intrinsic correlations and are also effective in accurate counterfactual estimation, they often lead to predictions with high variance or uncertainty estimates.
@@ 4.7 Time-invariant Causal Inference Methods
Time-invariant treatment effect estimation methods focus on quantifying the causal impact of a treatment on an outcome when both the treatment and outcome variables are observed at a single point in time. Whereas, a treatment is time-invariant or fixed when it occurs at one specific point of time and then does not change afterwards. We enlist some of the time-invariant causal inference methods below:
@@ 4.7.1 Causal-ARIMA
The Causal-ARIMA (C-ARIMA) is a potential outcome framework-based method to estimate causal effects from time series datasets (Menchetti et al., 2021). This method is suitable for datasets where it is not possible to separate the treated and controlled groups to estimate the effects of the applied treatment. In an observational study, if an intervention event occurs at a specific time point, the time series divides into two different time horizons, pre-intervention and post-intervention. Here the intervention is not applied at each time point, it occurred at a given time and draws the effect on the rest of the observations of the system. The C-ARIMA method is a three-step process: first, the relationship between the covariates and target variable is learned from the pre-intervention data using an ARIMA model. Then using learned relationships from pre-intervention data the counterfactual outcomes are predicted for post-intervention data points. Finally, the causal effect is measured from the factual observations and predicted counterfactual data. The C-ARIMA method works on the basis of five assumptions: single persistent intervention, non-anticipating potential outcomes, temporal no-interference, covariates-treatment independence, and Non-anticipating treatment. The C-ARIMA method can identify different types of causal effects without making any structural assumptions about the effect caused due to the intervention.
@@ 4.7.2 Difference in Difference
The Difference-in-Differences (DID) method (Lechner, 2011) is a statistical technique used in econometrics and social sciences to estimate the causal impact of a treatment, intervention, or policy by comparing changes in outcomes over time between a treatment group and a control group. DID is particularly useful when randomization of subjects into treatment and control groups is not possible or ethical, making it challenging to establish causality through traditional experimental methods. To perform causal inference using this method, start by selecting two groups: a treatment group that is exposed to the intervention or policy change and a control group that is not exposed. These groups should be similar in all relevant aspects except for the treatment or policy change. Next, data is collected on the outcome of interest for both groups, typically before and after the treatment or policy change. Finally, the causal effect is estimated by calculating the difference in the changes in outcomes in the controlled group and the treated groups. The DID estimate can be analyzed using regression techniques to control for potential confounding factors, such as demographics or other trends affecting the outcome. However, there are some limitations. The method requires adequate sample size to yield meaningful results and it can handle time-varying confounders. Overall, Difference-in-Differences is a powerful tool for causal inference when randomized controlled trials are not feasible or ethical. However, it relies heavily on the validity of the parallel trends assumption, and researchers must carefully consider potential sources of bias when designing and interpreting DID studies. Mathematically, the DID estimate can be expressed as follows:
@@ 4.7.3 Interrupted Time Series
Interrupted time series (ITS) analysis (McDowall et al., 1980) is a statistical method used for assessing the causal impact of an intervention, treatment, or policy change on a specific outcome by analyzing changes in a time series data set. Unlike the Difference-in-Differences (DID) method, which compares two groups (treatment and control) before and after an intervention, ITS focuses on a single group and tracks changes in that group’s outcomes over time, before and after the intervention. ITS requires defining a point in time when the intervention, treatment, or policy change occurred. This is referred to as the ”interruption point”. A statistical model is built to estimate the expected trajectory of the outcome in the absence of the intervention. Common models used for pre-intervention estimation include linear regression, autoregressive integrated moving average (ARIMA), or segmented regression models. To estimate the effect of the intervention, observed outcomes after the intervention are compared with the predicted outcomes from the model. The difference between the observed and predicted outcomes at each time point after the intervention represents the intervention effect.
@@ 4.7.4 Causal Impact
The Causal Impact method (Brodersen et al., 2015) is a statistical approach used for estimating the causal effect of an intervention, event, or treatment on a time series data set. It was developed by Google’s Research and Data Science teams and is implemented in the R package (http://google.github.io/CausalImpact/CausalImpact.html). This method is particularly useful for businesses and researchers to assess the impact of marketing campaigns, policy changes, or any other interventions on metrics such as sales, website traffic, or user engagement. The Causal Impact method builds a Bayesian structural time series (BSTS) model to represent the expected behavior of the time series data in the absence of the intervention. The BSTS model captures trends, seasonality, and other patterns in the pre-intervention data. The difference between the observed post-intervention data and the counterfactual prediction represents the causal impact of the intervention. This difference is used to estimate the treatment effect and assess its statistical significance. Causal Impact is particularly valuable when you want to assess the impact of a specific event or intervention on a time series metric while accounting for the underlying patterns and variations in the data. It offers a probabilistic and data-driven approach to causal inference, which can be beneficial for decision-making in various domains.
@@ 4.7.5 Regression Discontinuity Design
Table 8: Comparison of existing time-series causal inference methods and their relevance to Earth Science.
@@ 4.8 Spatiotemporal Causal Inference
4.8.1 Spatially Interrupted Time-Series (SITS)
(Zhang and Ning, 2023) investigated the spatiotemporal heterogeneities in the causal effects of mobility intervention policies during the COVID-19 outbreak using a spatially interrupted time-series (SITS) analysis. The study aimed to understand the spatiotemporal heterogeneity in the causal effectiveness of these policies, particularly at finer-grained spatial and temporal resolutions, and provide insights for policymakers on when and where to implement specific policies to control the spread of the virus. The SITS system involves the application of the Interrupted Time-Series (ITS) design combined with spatial analysis. The ITS design is a quasi-experimental design for policy evaluation that records and measures behavioral outcomes before and after implementing a policy intervention. The SITS analysis requires a long time series with multiple observations over time, especially before the treatment period, to establish a credible estimation of counterfactuals. It compares mobility outcomes over time within a single group of individuals or neighborhoods, rather than between a treatment group and a control group. This helps to control for potential biases arising from between-group differences. To incorporate spatial heterogeneities, the SITS analysis considers the spatial contexts of neighborhoods, including socioeconomic and built environment settings. The effects of the policy intervention, such as the level and slope changes, are moderated by these spatial contexts. The SITS model estimates the spatiotemporally varying causal effects of the policy interventions by incorporating these spatial moderating effects. The SITS analysis in this study used a timeseries dataset of neighborhood-based mobility levels extracted from a large set of mobile phone signaling data in Shenzhen, China. The dataset allowed for tracking the daily mobility trend at the neighborhood level before and after the COVID-19 outbreak. Two mobility control policies were focused on: the first-level response to public health emergencies (FLR) declared on January 23, 2020, and the lockdown of residential communities (COM) implemented on February 7, 2020. By revealing the disparities in the ability to alter normal activity patterns in response to NPIs, the study also provides insights into sociospatial equality and resilient neighborhood planning in the post-epidemic era. The study analyzed the effects of two mobility intervention policies: the first-level response (FLR) policy and the closed-off management (COM) policy. The spatial distributions of the decay rates of the two policies were different, with neighborhoods with larger decay rates primarily located in compact areas near city centers for the FLR and at the urban fringe for the COM. The study also found that the spatial variations in mobility change became increasingly small or homogenous over time.
@@ 4.9 Applications of Causal Inference in Earth Science
Applications of causal inference in Earth Science is constrained by the inability of validating the findings as observational data cannot be controlled, randomized or intervened on. Nevertheless, we enlist some of the recent inroads in causal inference for Earth Science domains. (Yang et al., 2022) applied an advanced data-driven causal inference method based on the convergent cross mapping (CCM) algorithm that provides valuable insights to identify thermal influence and feedback among the different climate regions in the contiguous U.S using the long-term (1901–2018) near surface air temperature observations. (Gao et al., 2022) studied some widely applied causal inference approaches to study the extent of the influence of different vegetation structures on net primary production (NPP) from both temporal and spatial perspective. Granger Causality Test (GCT), CCM, etc. are some of the used temporal causal models in this study. The authors conclude that in scenarios where it is difficult to infer causation from a temporal perspective, it may be feasible to use spatial causal inference there. (Wu et al., 2023) used a CCM method to study the the driving effects of various environmental factors such as air temperature, vegetation index, soil moisture, net surface radiation, precipitation, water vapor, etc. on land surface temperature in China based on the remote-sensing and reanalysis data from 2003–2018.
@@ 5 Datasets and Toolbox for Causality
Datasets play an important role when it comes to validating and analyzing the effectiveness of any new model or method. The ground truth knowledge of the dataset helps to approve or reject the result produced through any model. In this section, we provide reference to multiple synthetic, simulation and real-world datasets used for benchmarking causal discovery and inference methods for Earth Science domain. A summary of the publicly available toolboxes for causal study is given in Table 9.
5.1 Synthetic datasets
This type of data is generated using self-defined processes/equations by introducing different statistical and distribution features. So we know the structure of the data before applying the causal analysis method which is helpful to computer different evaluation matrices through ground truth.
Harvard Dataverse:
Six synthetic datasets are available at Harvard Dataverse101010 https://dataverse.harvard.edu/dataverse/basic_causal_structures_additive_noise, where these datasets represent different causal structures such as fork, v-structure, mediator, diamond, and seven variables, etc. The time series of each dataset is generated using a nonlinear function of cause variables, linear self-causation, and additive Gaussian noise. The number of time steps of all time series is set to 1000.
FLAIRS:
(Huang and Kleinberg, 2015) is a set of simulated time series datasets for causal analysis. This resource contains 22 simulated datasets, two datasets with a common cause and common effect, and 20 datasets each with a different random causal structure. All datasets contain 20 variables, 1000 time points, lags of [1,3] time units and all variables have continuous values.
FinanceCPT:
(Kleinberg, 2013) is a simulated financial time series dataset collection that contains 25 variables and 10 causal structures. For each causal structure two different observation time periods of 4000 time steps are generated, so this collection has 20 different datasets. The two observation periods of the same causal structure help to test any causal model that can find the same structure in both periods. These datasets have [20, 40] random relationships between variables with 1-3 time lags
PROMO dataset:
(Guyon et al., 2011) is a simple causal effect time series dataset of product sales and different promotions applied for each product. This dataset contains about 1000 promotion types applied on 100 different products. The daily sales value for each product is recorded for three years (i.e., 1095 days). So the daily sales and active promotions of each day for each product are available in the dataset. The goal is to predict the effect of each promotion on daily sales of each product. The ground truth causal relationship matrix is available in the dataset to evaluate the predicted result.
5.2 Realistic Simulated Data
In physical systems like the Earth system, is it not possible to perform randomized control experiments to collect data for studying causal properties. So the acceptable alternative is the computerized simulation of the physical systems which implements almost all dynamic processes to generate data resembling the actual system. Though these simulations are sometimes very time-consuming and expensive but help to produce data by changing different factors and are possible to generate counterfactual data.
North American Mesoscale:
(NAM) dataset (National Centers for Environmental Prediction, National Weather Service, NOAA, U.S. Department of Commerce, 2015) is generated by the National Centers for Environmental Prediction (NCEP) using the WRF Non Hydrostatic Mesoscale Model. This is a spatiotemporal dataset of 12km resolution covering the continental United States and the data is generated every 6 hours from 2012-01-01 00:00 to 2023-10-15 18:00. Different properties of Air Temperature, Geopotential Height, Humidity, Sea Level Pressure, Snow, Surface Pressure, Upper Air Temperature, and Upper Level Winds are available in this simulation.
FMRI benchmark:

### v2

@@ 4.5.1 Marginal Structural Models
Marginal Structural Models (MSMs) (Robins et al., 2000) comprise methods designed to estimate the effects of time-varying treatments while tackling the challenge of time-varying confoundedness. Since traditional methods like ordinary regression models may not properly handle time-varying treatments and confounders, MSMs were developed to address biases that can arise when using these traditional regression models to estimate causal effects in situations where treatments change over time. MSMs provide a framework to model and adjust for the dynamic nature of treatments, confounders, and their interdependencies using the IPTW weights. MSMs first employ IPTW to re-weight the data and emulate a hypothetical time-fixed treatment scenario. This mitigates confounding by making treated and untreated groups comparable. A weighted regression model is then employed to estimate the treatment effect, accounting for the dynamic confounding. MSMs rely on the assumptions of positivity and no unmeasured confounders. Though MSMs have become a cornerstone in time-varying causal inference in the fields of epidemiology (VanderWeele, 2009; Hernán et al., 2000), public health (Williamson and Ravani, 2017; VanderWeele et al., 2011) and social sciences(Bacak and Kennedy, 2015), they also have some limitations. Complex interactions between time-varying treatments and confounders, longitudinal missing data and information censoring can be challenging to model using MSM technique.
@@ 4.5.2 Convergent Cross Mapping
Convergent Cross Mapping (CCM) (Ye et al., 2015) is a nonlinear causal inference technique used to detect the presence of causal relationships in time-varying data. It is particularly useful when the relationship between variables is complex and non-linear. CCM is based on the concept of time delay embedding, where time series data is transformed into higher-dimensional space by embedding time-delayed copies of the series.
@@ 4.5.4 Deep Representation Learning based Models
In recent years, there has been growing interest in leveraging the powerful representation learning capabilities of deep learning models to tackle challenges in causal inference (Bengio et al., 2013; Koch et al., 2021). By encoding complex relationships in the data, deep learning based CI methods can learn a lower-dimensional representation of covariates, which can be used to predict counterfactual outcomes, treatment effect estimation and even control for confounding. Neural networks can also be used to estimate propensity scores, which represent the probability of treatment assignment given covariates. These propensity scores can then be used for matching, weighting, or as covariates in outcome models. For time series data, deep learning models can be used to handle the temporal dependencies and estimate causal effects over time (Moraffah et al., 2021). Recurrent Marginal Structural Networks (R-MSN) (Lim, 2018), Time-series deconfounder (Bica et al., 2020a) and Counterfactual Recurrent Network (CRN) (Bica et al., 2020b) are some of the deep learning based causal inference techniques that deploy g-methods to tackle time-varying confounding, however these techniques can only work for binary treatment effect estimation. Whereas, TCINet (Ali et al., 2023) and G-Net (Li et al., 2021b) can work for both binary and continuous treatment effect estimation, making them more suitable for Earth science domains.
@@ 4.6 Time-invariant Causal Inference Methods
Time-invariant treatment effect estimation methods focus on quantifying the impact of an intervention on the outcome when both the intervened and un-intervened outcome variables are observed at the same time. Whereas, the effect of time-invariant intervention is measured based on the difference in the outcomes before and after the intervention takes place. We enlist some of the time-invariant causal inference methods below:
@@ 4.6.1 Causal-ARIMA
The Causal-ARIMA (C-ARIMA) is a potential outcome framework-based method to estimate causal effects from time series datasets (Menchetti et al., 2021). This method is suitable for datasets where it is not possible to separate the treated and controlled groups to estimate the effects of the applied treatment. In an observational study, if an intervention event occurs at a specific time point, the time series divides into two different time horizons, pre-intervention and post-intervention. Here the intervention is not applied at each time point, it occurred at a given time and draws the effect on the rest of the observations of the system. The C-ARIMA method is a three-step process: first, the relationship between the covariates and target variable is learned from the pre-intervention data using an ARIMA model. Then using learned relationships from pre-intervention data the counterfactual outcomes are predicted for post-intervention data points. Finally, the causal effect is measured from the factual observations and predicted counterfactual data. The single persistent intervention, non-anticipating prospective outcomes, temporal no-interference, independence of covariates and treatment, and unforeseen treatment are the five underlying principles that underpin the C-ARIMA method’s operation. The C-ARIMA method can identify different types of causal effects without making any structural assumptions about the effect caused due to the intervention.
@@ 4.6.2 Difference in Difference
The Difference-in-Differences (DID) method (Lechner, 2011) is a statistical technique used in econometrics and social sciences, that considers a treatment group and a control group to assess the causal influence of a treatment, intervention, or policy by comparing changes in outcomes between two groups over time. DID is particularly useful when randomization of subjects into treatment and control groups is not possible or ethical, making it challenging to establish causality through traditional experimental methods. To perform causal inference using this method, start by selecting two groups: a treatment group that is exposed to the intervention or policy change and a control group that is not exposed. These groups should be similar in all relevant aspects except for the treatment or policy change. Next, data is collected on the outcome of interest for both groups, typically before and after the treatment or policy change. Finally, the causal effect is estimated by calculating the difference in the changes in outcomes in the controlled group and the treated groups. The DID estimate can be analyzed using regression techniques to control for potential confounding factors, such as demographics or other trends affecting the outcome. However, there are some limitations. The method requires adequate sample size to yield meaningful results and it can handle time-varying confounders. Overall, Difference-in-Differences is a powerful tool for causal inference when randomized controlled trials are not feasible or ethical. However, it relies heavily on the validity of the parallel trends assumption, and researchers must carefully consider potential sources of bias when designing and interpreting DID studies. The mathematical formula to quantify the effect of the applied treatment is given as:
@@ 4.6.3 Interrupted Time Series
Interrupted time series (ITS) analysis (McDowall et al., 1980) is a statistical method used for assessing the causal impact of an intervention, treatment, or policy change on a specific outcome by analyzing changes in a time series data set. Unlike the Difference-in-Differences (DID) method, which compares two groups (treatment and control) before and after an intervention, ITS focuses on a single group and tracks changes in that group’s outcomes over time, before and after the intervention. ITS requires defining a point in time when the intervention, treatment, or policy change occurred. This is referred to as the ”interruption point”. A statistical model is built to estimate the expected change direction of the outcome in the absence of the intervention. Common models used for pre-intervention estimation include linear regression, autoregressive integrated moving average (ARIMA), or segmented regression models. The intervention effect is defined as the variation between the expected and observed results at each time point following the intervention.
@@ 4.6.4 Causal Impact
The Causal Impact method (Brodersen et al., 2015) is a statistical approach for temporal data to estimate the causal effect of an intervention, event, or treatment. It was developed by Google’s Research and Data Science teams and is implemented in the R package  [^10] http://google.github.io/CausalImpact/CausalImpact.html. This method is particularly useful for businesses and researchers to assess the impact of marketing campaigns, policy changes, or any other interventions on metrics such as sales, website traffic, or user engagement. The Causal Impact method builds a Bayesian structural time series (BSTS) model to represent the expected behavior of the time series data in the absence of the intervention. The BSTS model captures trends, seasonality, and other patterns in the pre-intervention data. The difference between the observed post-intervention data and the counterfactual prediction represents the causal impact of the intervention. This difference is used to estimate the treatment effect and assess its statistical significance. Causal Impact is particularly valuable when you want to assess the impact of a specific event or intervention on a time series metric while accounting for the underlying patterns and variations in the data. It offers a probabilistic and data-driven approach to causal inference, which can be beneficial for decision-making in various domains.
@@ 4.6.5 Regression Discontinuity Design
Table 8: Comparison of existing causal inference methods for temporal data and their relevance to Earth science.
@@ 4.7.1 Propensity Score Based Spatiotemporal Causal Inference Methods
4.7.2 Spatially Interrupted Time-Series (SITS)
(Zhang and Ning, 2023) investigated the spatiotemporal heterogeneities in the causal effects of mobility intervention policies during the COVID-19 outbreak using a spatially interrupted time-series (SITS) analysis. The (Zhang and Ning, 2023) conducted a study to understand the spatiotemporal heterogeneity and the causal effectiveness of mobility intervention policies during COVID-19 outbreak, particularly at finer-grained spatial and temporal resolutions. Their research provided insights for policymakers on when and where to implement specific policies to control the spread of the virus. The SITS system involves the application of the Interrupted Time-Series (ITS) design combined with spatial analysis. The ITS design is a quasi-experimental design for policy evaluation that records and measures behavioral outcomes before and after implementing a policy intervention. The SITS analysis requires a long time series with multiple observations over time, especially before the treatment period, to establish a credible estimation of counterfactuals. It compares mobility outcomes over time within a single group of individuals or neighborhoods, rather than between a treatment group and a control group. This helps to control for potential biases arising from between-group differences. To incorporate spatial heterogeneities, the SITS analysis considers the spatial contexts of neighborhoods, including socioeconomic and built environment settings. The effects of the policy intervention, such as the level and slope changes, are moderated by these spatial contexts. The SITS model estimates the spatiotemporally varying causal effects of the policy interventions by incorporating these spatial moderating effects. The SITS analysis in this study used a time-series dataset of neighborhood-based mobility levels extracted from a large set of mobile phone signaling data in Shenzhen, China. The dataset allowed for tracking the daily mobility trend at the neighborhood level before and after the COVID-19 outbreak. The study analyzed the effects of two mobility intervention policies: the first-level response (FLR) policy and the closed-off management (COM) policy.
4.7.3 Spatio-Temporal Causal Inference Network (STCINet)
Most recently, (Ali et al., 2024) proposed a deep learning based causal inference technique by extending the potential outcome framework for spatiotemporal data. Their proposed method, STCINet, is a multi-input model with a U-Net based stream for counterfactual prediction and a latent factor model for handling time-varying confounders. STCINet works under the assumption of no unmeasured confounding and is able to estimate treatment effects of space-time-varying interventions. The goal is two handle two major challenges in performing spatiotemporal causal inference. One, to reduce bias caused by time-varying confounders. Two, to incorporate spatial interference into the outcome predictions. The authors evaluate their model on two diffusion based synthetic datasets, with and without spatial interference, while they demonstrate the performance of their method on real-world Arctic sea ice predictions by estimating the impact of atmospheric interventions on summer sea-ice forecasting in different sub-regions within the Arctic.
@@ 4.8 Applications of Causal Inference in Earth Science
In Earth science applications of causal inference are constrained by the inability to validate the findings as observational data cannot be controlled, randomized or intervened on. Nevertheless, we enlist some of the recent inroads in causal inference for Earth science domains. (Yang et al., 2022) applied an advanced data-driven causal inference method based on the convergent cross mapping (CCM) algorithm that provides valuable insights to identify thermal influence and feedback among the different climate regions in the contiguous U.S using the long-term (1901–2018) near surface air temperature observations. (Gao et al., 2022) studied some widely applied causal inference approaches to study the extent of the influence of different vegetation structures on net primary production (NPP) from both temporal and spatial perspective. Granger Causality Test (GCT), CCM, etc. are some of the used temporal causal models in this study. The authors conclude that in scenarios where it is difficult to infer causation from a temporal perspective, it may be feasible to use spatial causal inference there. (Wu et al., 2023) used a CCM method to study the driving effects of various environmental factors such as air temperature, vegetation index, soil moisture, net surface radiation, precipitation, water vapor, etc. on land surface temperature in China based on the remote-sensing and reanalysis data from 2003–2018.
@@ 5 Datasets and Toolbox for Causality
Datasets play a vital role in validating and analyzing the effectiveness of any new model or method. The ground truth knowledge of the dataset helps to approve or reject the result produced through any model. In this section, we provide reference to multiple synthetic, simulation and real-world datasets that can be utilized in causal discovery and inference methods benchmarking in different science domains. A summary of the publicly available toolboxes for causal study is given in Table 11.
Table 9: Synthetic, simulated, and real world datasets used for causal analysis.
Dataset Description Data Causal Dataset
Type Category Type
Harvard Dataverse  [^11] https://dataverse.harvard.edu/dataverse/basic_causal_structures_additive_noise Contains six synthetic datasets representing different causal structures. The time series datasets are generated using Time Causal Synthetic
a nonlinear function of cause variables, linear self-causation and additive Gaussian noise. series Discovery
FLAIRS This resource contains 22 simulated time series datasets. All datasets contain 20 continuous variables and Time Causal Synthetic
(Huang and Kleinberg, 2015) 1000 time points with the lag of 1 and 3 time units. series Discovery
FinanceCPT This simulated financial dataset collection contains 20 datasets with 25 variables and 10 causal structures. So two Time Causal Synthetic
(Kleinberg, 2013) datasets for the same causal structure. These datasets have [20, 40] random relationships between variables with series Discovery
1-3 time lags and 4000 time steps.
PROMO This is a dataset of product sales and different promotions for each product. The daily sales value for each Time Causal Synthetic
(Guyon et al., 2011) product is recorded for three years with ground truth graph where about 1000 promotions applied on series Discovery
100 different products.
Diffusion Data This dataset contains 4000 samples of diffusion-based spatiotemporal images. The dataset contains 3 variables Spatio- Causal Synthetic
(Ali et al., 2024) including treatment, time-varying confounder, and (factual and counterfactual) outcomes. temporal Inference
North American Mesoscale Generated by the National Centers for Environmental Prediction (NCEP) using the WRF Non-Hydrostatic Mesoscale Spatio- Causal Realistic
(NAM) Model. This is a spatiotemporal dataset of 12km resolution covering the continental United States and the data frequency temporal Discovery Simulated
(NCEP, NWS, NOAA, 2015) is every 6 hours from 2012-01-01 00:00 to 2023-10-15 18:00. Different properties of Air Temperature, Geopotential
Height, Humidity, Sea Level Pressure, Snow, Surface Pressure and Upper Level Winds are available in this simulation.
FMRI benchmark Functional Magnetic Resonance Imaging (FMRI) dataset contains 28 distinct simulations for various brain networks. Time Causal Realistic